40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6027     Entry time: Mon Nov 28 16:51:57 2011
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: modulation frequency reset 

I reset the modulation frequency to 11065910 Hz (#5530). It had been at 11065399 Hz probably since the power shut down.

ELOG V3.1.3-